OpenStudy (anonymous):

two forces of 509 N and 232 N at a point .the resultant force is 585 N find the angle between the forces. help me plz

OpenStudy (ybarrap):

is this what's being asked?  The wording is a bit ambiguous

OpenStudy (anonymous):

ye

OpenStudy (anonymous):

can u help me plz

OpenStudy (ybarrap):

You can try this. Create a coordinate system where the resultant force is on the "y-axis" and the force perpendicular to the resultant force is the "x-axis":  The the sum along the y-axis will be R and the sum along the x-axis will be zero: $$ F_1\cos\theta_1+F_2\cos\theta_2=R\\ F_1\sin\theta_1-F_2\sin\theta_2=0 $$ Two equations, two unknowns. Solvable. Make sense?